ANOVA df SS Regression 1 882 Residual 20 4000 Total 21 4882 Coefficients Standard Error t...

ANOVA

df

SS

Regression

1

882

Residual

20

4000

Total

21

4882

Coefficients

Standard Error

t Stat

Intercept

5.00

3.56

Variable x

6.30

3.00

Use the ANOVA table that was provided in question 7 and Perform an F test and determine whether x and y are related. Use α = .05

Answer Options:

Since the test statistic F = 3.45 < 4.35 ,Fail to reject HO

Since the test statistic F = .45< 3.45, Fail to reject HO

Since the test statistic F = 4.41 > 4.35, reject H0.

Since the test statistic F = 4.41>4.05, reject HO.
Answer #1

We have F test statistic value is

F=\frac{\frac{882}{1}}{\frac{4000}{20}}=4.41

F critical value is 4.35.........................by using =FINV(0.05,1,20)

F test statistic value >F critical value

i.e. 4.41 > 4.35  

therefore we reject H0.

Since the test statistic F = 4.41 > 4.35, reject H0.